OpenStudy (anonymous):

i really need help with this, Find the percent error or the percent decrease: Glen is looking into buying a piece of land. He estimates that the land measures 100yd by 850yd. The land is actually 140yd by 940yd. What is the percent error in the estimated area of the land?

OpenStudy (anonymous):

estimate is \[100\times 850=85,000\] actual area is \[140\times 940=131,600\]

OpenStudy (anonymous):

I am not sure, .0987... in my opinion

OpenStudy (anonymous):

\[131,600-85,000=46,600\] a rather large error if i have my numbers right

OpenStudy (anonymous):

Righ! my answer is wrong

OpenStudy (anonymous):

as a percent it would be \[\frac{466,00}{131,600}=35.4\%\] unless i messed up

OpenStudy (anonymous):

Thank you soooo much